OpenStudy (needhelpp101):

~*~``BECOME FAN && MEDAL && TESTIMONIAL``~*~ Use the discriminant to find the number of real solutions of the equation 3x^2 – 5x + 4 = 0. Show your work.

OpenStudy (*poker-face*):

do you know how to find discriminant?

OpenStudy (needhelpp101):

nope :(

OpenStudy (needhelpp101):

This equation is of form ax^2+bx+c=0 a = 3 b = -5 c = 4 x= [-b+/-sq. (b^2-4ac)]/2a] x= [5 +/-sq. (-5^2-4(3)(4)]/(2)(3) Discriminant is b^2-4ac = -23

OpenStudy (needhelpp101):

@*Poker-Face* ???

OpenStudy (*poker-face*):

yes correct the discriminant is -23

OpenStudy (*poker-face*):

whenever the discriminant is negative it implies that the equation has no real roots

OpenStudy (needhelpp101):

Okay thx :)

OpenStudy (*poker-face*):

np (: